PLEASE HELP ME..!
Find the value of x in the matrix equation.

{{3x+2y, -6x-3y}} = {{6,-6}}

Answer by stanbon(75887)

Find the value of x in the matrix equation.
{{3x+2y, -6x-3y}} = {{6,-6}}
-----------------------
3x+2y = 6
-6x-3y = -6
----------------
Multiply the 1st equation by 2:
6x+4y = 12
-6x-3y = -6
----
Add and solve for "y":
y = 6
-------------
Solve for "x":
3x+2y = 6
3x+2*6 = 6
3x = -6
x = -2
=============
Cheers,
Stan H.
=============